OpenStudy (anonymous):

write a possible explicit rule for the nth term of each sequence 0.8 , 1.6 , 3.2 , 6.4 , 12.8

hartnn (hartnn):

do you find any pattern in that sequence ?

OpenStudy (tkhunny):

Have you considered successive ratios? 1.6 / 0.8 = 2 3.2 / 1.6 = ?? Hunt around a little.

OpenStudy (anonymous):

no , i dont get how to do this one at aall . :0

OpenStudy (freckles):

I think the detective tkunny was on to something there.

hartnn (hartnn):

there's a pattern! if you see each term is 2 times the previous term!

OpenStudy (freckles):

the hint was ratio

OpenStudy (anonymous):

ohkay i see what you mean finally about the term is 2 times the previous term , but now what ..

OpenStudy (anonymous):

is that the answer? loll this is so wackkkk to me.

OpenStudy (tkhunny):

Write an explicit rule. You do not have "the answer" until you write the rule. Write!

OpenStudy (tkhunny):

The first rule would be never to say that again - or think it. Just let it go. What's the first element in the sequence? How can we get to the second element from the first?

OpenStudy (anonymous):

by multiplying by 2 every time you get an answer. ?

hartnn (hartnn):

common ratio(r) is the ratio between consecutive terms, here r =2 1st term = a1 = 0.8 use the general formula \(a_n = a_1 r^{n-1} \) just plug in values!

OpenStudy (anonymous):

so its an=0.8 2n-1?

hartnn (hartnn):

yes, \(a_n = 0.8 (2)^{n-1} = (0.8/2) (2^n) =0.4 (2)^n \)

OpenStudy (anonymous):

an= 0.8(2) n-1 is the main answer tho right? thats all i have to put? or the other stuff to?

OpenStudy (reemii):

with a little hat : 0.8 * (2^(n-1))

OpenStudy (anonymous):

?

hartnn (hartnn):

\(a_n = 0.8 (2)^{n-1}\) is the correct answer and \(a_n = 0.4 (2)^{n}\) is the simplified form you can put either of them

OpenStudy (reemii):

he's right. my intervention is to point out that : 0.8 (2) n-1 is not going to be understood by a "robot" as \(0.8 (2^{n-1})\). if you write text, you have to use the '^' (hat) which means "to the power".